9. In a production process, one fifth of the items fabricated are defective. Ten items from the production line are randomly selected and inspected. Let X be the number of defective articles in this sample. If the distribution of X is Bin(n,p), what are the values of n and p? 10. of the random variable X is given. What is the mean of the distribution? 11. What is of the distribution? 12, X is a random normal normal variable, with mean and variance ơ. The-standardized form of X is Z = (X-Do: What are the mean and variance (u, σ), respectively, of Z? 13. Suppóse X N(5,3). What is P(X s 8) in terms of the standard normal variable Z? 14. Suppose X N(5,3) What is the value of PX2)? 15. Suppose X is normally distributed with mean 5 and standard deviation 0.4. Using the standard transformation Z X Vo we find Ps X)PIZ s 1.3). What is the value of X? X 3.52 16. Suppose X is normally distributed with mean 5. If PK <3) 0.2 what is the standard distribution of X? σ: 3-5 T: 2.38075 238
Number 12
Answer #1

A normal distribution with mean=0 and standard deviation =1 is called standard normal distribution. (Z distribution)

So mean=0 and

variance=sd^2

=1^2=1